OpenStudy (anonymous):

Please help me... I will be ur fan....plzzzzz :**** Part 1: Create a scenario for an arithmetic sequence. For example, Jasmine practices the piano for ______ minutes on Monday. Every day she ___________ her practice time by _________. If she continues this pattern, how many minutes will she practice on the 7th day? Use your scenario to write the function for the 7th term in your sequence using sequence notation. Part 2: Create a scenario for a geometric sequence. For example, Anthony goes to the gym for ____ mi. Use your scenario to write the formula for the 5th term in your sequence using sequence notation. Part 3: Use your scenario from part 2 to write a question that will lead to using the geometric series formula. Use the formula to solve for Sn in your scenario

OpenStudy (study_together):

So should we write another or solve the same blank ones above?

OpenStudy (moazzam07):

i have a data

OpenStudy (moazzam07):

this should be helpful

OpenStudy (anonymous):

I think we need to right another one...

OpenStudy (study_together):

ok

OpenStudy (anonymous):

yeah.....

OpenStudy (anonymous):

so can you help me?

OpenStudy (moazzam07):

These are the 5 numbers that you should list for part 2: Mon - 20 minutes Tue - 22 minutes Wed - 24.2 minutes Thu - 26.62 minutes Fri - 29.282 minutes

OpenStudy (kash_thesmartguy):

You do it. Fill in the blanks first and then turn it into an equation!

OpenStudy (moazzam07):

i did this problem before i just dont have the data for Part 1 anymore

OpenStudy (anonymous):

but I need another story..

OpenStudy (study_together):

Try this one Blake goes running 30 mins on monday and continue on...

OpenStudy (kash_thesmartguy):

\(\color{blue}{\text{Originally Posted by}}\) @Kash_TheSmartGuy You do it. Fill in the blanks first and then turn it into an equation! \(\color{blue}{\text{End of Quote}}\) For part one. It will help you

OpenStudy (study_together):

@plzzhelpme

OpenStudy (anonymous):

Do u have the data for part 2?? @moazzam07

OpenStudy (anonymous):

and I'll give it try @Study_together

OpenStudy (study_together):

okay

OpenStudy (anonymous):

and @Kash_TheSmartGuy

OpenStudy (moazzam07):

@plzzhelpme that is the Data for Part 2

OpenStudy (kash_thesmartguy):

Create a scenario for an arithmetic sequence. For example, Jasmine practices the piano for 40 minutes on Monday. Every day she increases her practice time by 2 minutes. If she continues this pattern, how many minutes will she practice on the 7th day? Use your scenario to write the function for the 7th term in your sequence using sequence notation. 40+2x=x would result as the equation. This was just an example, do not copy it**

OpenStudy (anonymous):

I Understand a bit more.. and i know its an example...I'm gonna try to use the brian goes running story... but if I need any help I'll ask again... :***

OpenStudy (anonymous):

Blake*

OpenStudy (anonymous):

@Kash_TheSmartGuy for the equation u gave in the example.. did u mean 40+2^x=x or 40+2x=x??? and its asking for the 7 day so where do i substitute that?

OpenStudy (moazzam07):

These are the 5 numbers that you should list for part 2: Mon - 20 minutes Tue - 22 minutes Wed - 24.2 minutes Thu - 26.62 minutes Fri - 29.282 minutes

OpenStudy (anonymous):

Is this correct.... 40+2^7=x...?????

OpenStudy (anonymous):

This is what I've got so far.. is it correct????.... Blake goes running for 40 minutes on Sunday. Every day he increases his running time by 2 minutes. If he continues this pattern, how many minutes will he run on the 7th day? 40+2^x=x After 7 days: x=40+2^7 x=168 minutes what do I do for this question: Use your scenario to write the function for the 7th term in your sequence using sequence notation. @moazzam07 @Kash_TheSmartGuy @Study_together

OpenStudy (moazzam07):

i can help you later i gtg but ask @dtan5457

OpenStudy (anonymous):

ummm its ok...thx tho

OpenStudy (moazzam07):

ok just make an equation using the Data you have from the Scenario

OpenStudy (moazzam07):

are you good now. A?

OpenStudy (anonymous):

ya....ig .....thanks guys:)... :*